# Unveiling the Learning Mind of Language Models: A Cognitive Framework and Empirical Study

### v1

@@ Abstract
Large language models (LLMs) have shown impressive capabilities across tasks such as mathematics, coding, and reasoning, yet their learning ability, which is crucial for adapting to dynamic environments and acquiring new knowledge, remains underexplored. In this work, we address this gap by introducing a framework inspired by cognitive psychology and education. Specifically, we decompose general learning ability into three distinct, complementary dimensions: Learning from Instructor (acquiring knowledge via explicit guidance), Learning from Concept (internalizing abstract structures and generalizing to new contexts), and Learning from Experience (adapting through accumulated exploration and feedback). We conduct a comprehensive empirical study across the three learning dimensions and identify several insightful findings, such as (i) interaction improves learning; (ii) conceptual understanding is scale-emergent and benefits larger models; and (iii) LLMs are effective few-shot learners but not many-shot learners. Based on our framework and empirical findings, we introduce a benchmark that provides a unified and realistic evaluation of LLMs’ general learning abilities across three learning cognition dimensions. It enables diagnostic insights and supports evaluation and development of more adaptive and human-like models.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Abdin et al., 2024; Yang et al., 2024) have demonstrated impressive capabilities across diverse tasks such as mathematics (Zhang et al., 2024b, a; Yue et al., 2024), coding (Nam et al., 2024; Chew et al., 2023; Kim et al., 2024a), and reasoning (Hao et al., 2023; Yuan et al., 2024a; Zheng et al., 2023). Despite these achievements, as LLM-driven systems increasingly integrate into societal roles, serving as AI companions (Xu et al., 2024b, b; Chen et al., 2024; Ploderer et al., 2025) or AI employees (Brachman et al., 2025; McDuff et al., 2025; Singhal et al., 2025; Dillion et al., 2025), they must possess robust learning capabilities. Effective learning is essential for AI agents to adapt dynamically to diverse and changing environments, continuously acquire new knowledge, and autonomously respond to novel contexts. However, research on the learning ability of LLMs remains scarce, with little work systematically investigating how well LLMs can acquire and generalize new knowledge across tasks.
To address this gap, we aim to systematically investigate the learning ability of LLMs. Drawing insights from cognitive science and educational theory (Fodor, 1983; Kolb, 1984; Clark et al., 2012), we first analogize from human learning processes to identify three fundamental dimensions of learning: (1) learning from instructor, where the model acquires knowledge through guided interaction (Cole et al., 2012; Sheffield et al., 2018); (2) learning from concept, where the model internalizes structured abstractions and generalizes them to downstream tasks (Minsky, 1961; Beckmann et al., 2023); and (3) learning from experience, where the model adapts based on accumulated trajectories or exploration-feedback (Kolb, 1984; Zhao et al., 2024). For each dimension, we design targeted experimental paradigms to operationalize the corresponding learning mechanism. In Learning from Instructor, we simulate tutor-learner settings with and without interactive clarification, demonstrating that interaction within instructor and learner consistently improves model’s leaning ability. In Learning from Concept, we evaluate the impact of injecting abstract conceptual knowledge in competitive environments (i.e., TextArena (Guertler et al., 2025)), showing that (i) conceptual understanding is scale-emergent, and (ii) injecting structured domain knowledge can provide a tangible advantage, if the model is sufficiently capable of internalizing it. In Learning from Experience, it is a crucial capability for adapting to novel environments and acquiring new knowledge autonomously. While they are few-shot learners, they struggle in many-shot settings due to the challenges of long-context integration. This highlights the importance of a unified benchmark that can evaluate LLMs’ general learning abilities across cognitive dimensions.
@@ 2 Related Work
LLMs have been extensively evaluated on tasks involving linguistic competence (Srivastava et al., 2022; Chen et al., 2023; Kim et al., 2024b), factual recall (Geva et al., 2023; Wang et al., 2023), reasoning (Hao et al., 2023; Yuan et al., 2024a; Zheng et al., 2023), instruction following (Hu et al., 2024c, d, b), and multitask generalization (Muennighoff et al., 2023; Cohen et al., 2025; Yin et al., 2024), with benchmarks such as MMLU (Hendrycks et al., 2020b), BIG-Bench (Aarohi Srivastava, 2022), and HELM (Liang et al., 2023) highlighting broad capabilities and emergent scaling trends. However, most evaluations focus on static zero- or few-shot performance, offering limited insight into how models learn or generalize from experience. To address this, recent studies have begun shifting toward more dynamic formulations that probe the learning behavior of LLMs. (Dong et al., 2022) categorize in-context learning (ICL) as inference-time adaptation and identify key factors for success. (Min et al., 2022) show that ICL performance depends more on input format than label correctness. (Agarwal et al., 2024) find that many-shot ICL yields diminishing returns as context length grows. (Schick et al., 2023) introduce Toolformer, enabling models to self-supervise API use. (Yuan et al., 2024b) develop models that improve via internally generated feedback. Nevertheless, most such studies are piecemeal, lacking a unified theoretical lens or systematic analysis for comparing learning processes across tasks. Our work addresses this gap by introducing a cognitively grounded decomposition to systematically analyze learning ability.
@@
Recent advances in large language models have demonstrated remarkable capabilities in adapting and generalizing knowledge beyond traditional static evaluations (Brown et al., 2020; Wei et al., 2022; Tan et al., 2024). Cognitive psychology and educational theory have long emphasized that effective learning involves not only direct instruction and guided feedback (Clark et al., 2012; Cole et al., 2012), but also the ability to abstract conceptual structures and adapt from experience (Kolb, 1984; Minsky, 1961; Fodor, 1983). Motivated by these insights, we adopt a tripartite perspective, learning from instructor, concept, and experience, that mirrors rapid-instruction-to-implementation learning (RITL) in humans (Sheffield et al., 2018), classical rule-based accounts of cognition (Minsky, 1961; Fodor, 1983), and experiential learning theory (Kolb, 1984). While contemporary LLM benchmarks extensively measure task-specific performance (Aarohi Srivastava, 2022; Liang et al., 2023; Hendrycks et al., 2020b; Meva & Kukadiya, 2025), they rarely probe how models internalize instructions, extract explicit rules, or accumulate and reuse episodic experience, capacities increasingly highlighted by self-evolving or self-refinement agents (Gao et al., 2024; Hu et al., 2024a). This gap underscores the need for a cognitively informed evaluation framework that systematically characterizes the general learning abilities of LLMs across these complementary dimensions.
@@ 3 A Cognitive Framework for Analyzing Learning Abilities in LLMs
To systematically investigate the general learning capabilities of LLMs, we propose a cognitively grounded framework comprising three paradigms: Learning from Instructor (LfI), Learning from Concept (LfC), and Learning from Experience (LfE). LfI captures learning via explicit guidance, supported by evidence that instruction accelerates task acquisition and reduces cognitive load (Cole et al., 2012; Sheffield et al., 2018; Clark et al., 2012; Siregar, 2021). LfC reflects structured abstraction from static concepts, grounded in symbolic cognitive theories (Minsky, 1961; Fodor, 1983; Hu et al., 2024a) despite critiques of their rigidity (Beckmann et al., 2023). LfE models adaptation through feedback and interaction, consistent with experiential learning theory (Kolb, 1984; Huerta-Wong & Schoech, 2010) and recent work showing LLMs improve via self-refinement (Zhao et al., 2024; Gao et al., 2024).
@@ 4 Learning from Instructor (LfI)
We investigate Learning from Instructor (LfI), where models acquire task knowledge via structured guidance, such as demonstrations, explanations, or feedback. We evaluate this setting into two dimensions: (1) Passive Consumption vs. Interactive Clarification, and (2) Scaling Learner.
@@
To examine the impact of interactivity in instruction-based learning, we adopt the MagpieMath (Xu et al., 2024a) dataset, a high-quality math dataset generated autoregressively by Qwen2.5-Math-72B. An instructor model is first trained on this dataset and subsequently used to teach a separate learner model under two learning paradigms: (1) Passive Consumption, in which the learner receives only direct solutions from the instructor without any further interaction, and (2) Interactive Clarification, where the learner is allowed to ask clarification questions following each instructor response, and the instructor provides targeted, follow-up explanations. In this setting, the learner is restricted to a single clarification question per sample to ensure consistency across training examples. Both the initial answers and the clarification responses are aggregated as supervision for the learner. We experiment with four representative LLM families: LLaMA3.1-8B, Qwen2.5-7B, Mistral-7B, and Phi-3-mini—each used as both instructor and learner to test robustness across pairings. Learners are trained on outputs from matched instructors within each setting. Evaluation is conducted on eight diverse mathematical benchmarks: GSM8K (Cobbe et al., 2021), SVAMP (Patel et al., 2021), MATH (Hendrycks et al., 2021b), NumGLUE (Mishra et al., 2022), SimulEq (Koncel-Kedziorski et al., 2016), AQuA (Ling et al., 2017), SAT (Zhong et al., 2023), and MMLU-Math (Hendrycks et al., 2020a), covering competencies from basic arithmetic to multi-step symbolic reasoning and standardized test preparation. Implementation and generation details are provided in Appendix A.3.
@@
As shown in Figure 2, learners trained under the Interactive Clarification paradigm consistently outperform those trained via Passive Consumption across all eight evaluation benchmarks. This performance gain highlights that LLMs are capable of leveraging interactive feedback to improve task understanding, resembling human-like active learning behaviors. Notably, the gains vary across model families. Mid-sized models such as Mistral-7B and Qwen2.5-7B show substantial improvements, suggesting a strong capacity to benefit from additional instructional signals. In contrast, the smallest model, Phi-3-mini, shows only marginal improvement, indicating limited ability to engage in or benefit from interactive clarification, which suggest that active learning capabilities are not uniform across models and may depend on model capacity or architecture.
Figure 3: Scaling learner models (Qwen2.5 variants) with a fixed Qwen2.5-72B instructor. Larger learner learn more effectively, and interactive clarification further improves outcomes.
@@ 5 Learning from Concept (LfC)
Table 1: Accuracy on six tasks with and without concept input (baseline / +concept). Higher scores indicate better sub-goal completion or reasoning accuracy. Darker cell colors denote larger improvements; white means no change.
Qwen 2.5 7b 14b 32b 72b
[…]
@@ 5.1 Structured Knowledge Injection in Competitive Environments
We evaluate whether conceptual knowledge improves strategic performance in multi-agent settings using TextArena (Guertler et al., 2025), a suite of competitive environments with symbolic rules and multi-turn dynamics. Each game involves two players, Player-0 and Player-1. We fix Player-0 as Qwen2.5-32B and vary Player-1 across four scales: Qwen2.5-1.5B, 7B, 14B, and 32B. Prior to gameplay, Player-1 either receives no guidance (without Concept) or is given natural language descriptions of game rules and strategies (with Concept), generated by Qwen2.5-32B. For each environment and model pair, we run 20 matches and report Player-1’s win rate averaged over these games. Environments include Checkers (CH), Poker (PK), Stratego (ST), Tic Tac Toe (TT), Truth and Deception (TD), and Ultimate Tic Tac Toe (UTT). Performance is measured as Player-1’s win rate against the fixed opponent. Results are shown in Figure 4. Implementation details are provided in Appendix A.4.
@@ 5.2 Conceptual Generalization in Logic and Planning Tasks
To evaluate rule-based generalization, we include six tasks centered on explicit symbolic or logical structures. LogicGrid and NaLogic feature structured and narrative logic puzzles (Mandell, 1986; Dudeney, 1907), testing deduction under constraint. Plan (Vallati et al., 2015) translates classical symbolic planning domains (e.g., Gripper, Barman) into natural language settings to assess structured action reasoning. Additionally, tasks from AlfWorld (Shridhar et al., 2021), ScienceWorld (Wang et al., 2022a), and BabyAI (Chevalier-Boisvert et al., 2019), models must complete goal-oriented tasks requiring commonsense reasoning, procedural knowledge, and spatial navigation. For the concept-injected setting, we use Qwen2.5-72B to generate static conceptual hints, which are provided as auxiliary input to smaller models during evaluation. Results are presented in Table 1 and Figure 4. We use subgoal-based annotations to enable fine-grained progress tracking. Implementation details are provided in Appendix A.4.

### v3

@@ Abstract
Large language models (LLMs) have shown impressive capabilities across tasks such as mathematics, coding, and reasoning, yet their learning ability, which is crucial for adapting to dynamic environments and acquiring new knowledge, remains underexplored. In this work, we address this gap by introducing a framework inspired by cognitive psychology and education. Specifically, we decompose general learning ability into three distinct, complementary dimensions: Learning from Instructor (acquiring knowledge via explicit guidance), Learning from Concept (internalizing abstract structures and generalizing to new contexts), and Learning from Experience (adapting through accumulated exploration and feedback). We conduct a comprehensive empirical study across the three learning dimensions and identify several insightful findings, such as (i) interaction improves learning; (ii) conceptual understanding is scale-emergent and benefits larger models; and (iii) LLMs are effective few-shot learners but not many-shot learners. Based on our framework and empirical findings, we introduce a benchmark that provides a unified and realistic evaluation of LLMs’ general learning abilities across three learning cognition dimensions. It enables diagnostic insights and supports evaluation and development of more adaptive and human-like models. The code is available here  [^1] https://aka.ms/learnarena.
@@ 1 Introduction
Large language models (LLMs) (achiam2023gpt; abdin2024phi; yang2024qwen2; lian2024recai) have demonstrated impressive capabilities across diverse tasks such as mathematics (zhang2024mathverse; zhang2024mario; yue2024mathvc), coding (nam2024using; chew2023llm; kim2024exploring), and reasoning (Hao2023ReasoningWL; Yuan2024AdvancingLR; Zheng2023JudgingLW; liuyue_efficient_reasoning). Despite these achievements, as LLM-driven systems increasingly integrate into societal roles, serving as AI companions (xu2024can; chen2024learning; ploderer2025maintaining) or AI employees (brachman2025current; mcduff2025towards; singhal2025toward; dillion2025ai; huang2025recommender), they must possess robust learning capabilities. Effective learning is essential for AI agents to adapt dynamically to diverse and changing environments, continuously acquire new knowledge, and autonomously respond to novel contexts. However, research on the learning ability of LLMs remains scarce, with little work systematically investigating how well LLMs can acquire and generalize new knowledge across tasks.
To address this gap, we aim to systematically investigate the learning ability of LLMs. Drawing insights from cognitive science and educational theory fodor1983modularity; kolb1984experiential; clark2012putting, we first analogize from human learning processes to identify three fundamental dimensions of learning: (1) learning from instructor, where the model acquires knowledge through guided interaction cole2012rapid; sheffield2018rapid; (2) learning from concept, where the model internalizes structured abstractions and generalizes them to downstream tasks minsky1961steps; Beckmann2023AnAT; and (3) learning from experience, where the model adapts based on accumulated trajectories or exploration-feedback kolb1984experiential; zhao2024expel. For each dimension, we design targeted experimental paradigms to operationalize the corresponding learning mechanism. In Learning from Instructor, we simulate tutor-learner settings with and without interactive clarification, demonstrating that interaction within instructor and learner consistently improves model’s learning ability. In Learning from Concept, we evaluate the impact of injecting abstract conceptual knowledge in competitive environments (i.e., TextArena (guertler2025textarena)), showing that (i) conceptual understanding is scale-emergent, and (ii) injecting structured domain knowledge can provide a tangible advantage, if the model is sufficiently capable of internalizing it. In Learning from Experience, it is a crucial capability for adapting to novel environments and acquiring new knowledge autonomously. While they are few-shot learners, they struggle in many-shot settings due to the challenges of long-context integration. This highlights the importance of a unified benchmark that can evaluate LLMs’ general learning abilities across cognitive dimensions.
@@ 2 Related Work
LLMs have been extensively evaluated on tasks involving linguistic competence srivastava2022beyond; Chen2023InternVS; Kim2024HealthLLMLL, factual recall (Geva2023DissectingRO; Wang2023SurveyOF), reasoning (liuyue_GuardReasoner; liuyue_GuardReasoner_VL; Yuan2024AdvancingLR; Zheng2023JudgingLW; zhang2025gvpo; zhang2025rspo), instruction following hu2024rethinking; lei2024aligning; hu2024language; lei2024recexplainer; hu2024explaining, and multitask generalization Muennighoff2023CrosslingualGT; Cohen2025DFPEAD; Yin2024MultitaskBasedEO; hu2025population, with benchmarks such as MMLU hendryckstest2021, BIG-Bench Srivastava2022BeyondTI, and HELM Liang2023HolisticEO highlighting broad capabilities and emergent scaling trends. However, most evaluations focus on static zero- or few-shot performance, offering limited insight into how models learn or generalize from experience. To address this, recent studies have begun shifting toward more dynamic formulations that probe the learning behavior of LLMs. dong2022survey categorize in-context learning (ICL) as inference-time adaptation and identify key factors for success. Min2022RethinkingTR show that ICL performance depends more on input format than label correctness. Agarwal2024ManyShotIL find that many-shot ICL yields diminishing returns as context length grows. Schick2023ToolformerLM introduce Toolformer, enabling models to self-supervise API use. Yuan2024SelfRewardingLM develop models that improve via internally generated feedback. Nevertheless, most such studies are piecemeal, lacking a unified theoretical lens or systematic analysis for comparing learning processes across tasks. Our work addresses this gap by introducing a cognitively grounded decomposition to systematically analyze learning ability.
@@
Recent advances in large language models have demonstrated remarkable capabilities in adapting and generalizing knowledge beyond traditional static evaluations brown2020language; wei2022emergent; tan-etal-2024-democratizing. Cognitive psychology and educational theory have long emphasized that effective learning involves not only direct instruction and guided feedback clark2012putting; cole2012rapid, but also the ability to abstract conceptual structures and adapt from experience kolb1984experiential; minsky1961steps; fodor1983modularity. Motivated by these insights, we adopt a tripartite perspective, learning from instructor, concept, and experience, that mirrors rapid-instruction-to-implementation learning (RITL) in humans sheffield2018rapid, classical rule-based accounts of cognition minsky1961steps; fodor1983modularity, and experiential learning theory kolb1984experiential. While contemporary LLM benchmarks extensively measure task-specific performance Srivastava2022BeyondTI; Liang2023HolisticEO; hendryckstest2021; meva2025performance, they rarely probe how models internalize instructions, extract explicit rules, or accumulate and reuse episodic experience, capacities increasingly highlighted by self-evolving or self-refinement agents gao2024self; hu2024case. This gap underscores the need for a cognitively informed evaluation framework that systematically characterizes the general learning abilities of LLMs across these complementary dimensions.
@@ 3 A Cognitive Framework for Analyzing Learning Abilities in LLMs
To systematically investigate the general learning capabilities of LLMs, we propose a cognitively grounded framework comprising three paradigms: Learning from Instructor (LfI), Learning from Concept (LfC), and Learning from Experience (LfE). LfI captures learning via explicit guidance, supported by evidence that instruction accelerates task acquisition and reduces cognitive load cole2012rapid; sheffield2018rapid; clark2012putting; siregar2021explicit. LfC reflects structured abstraction from static concepts, grounded in symbolic cognitive theories minsky1961steps; fodor1983modularity; hu2024case despite critiques of their rigidity Beckmann2023AnAT. LfE models adaptation through feedback and interaction, consistent with experiential learning theory kolb1984experiential; huerta2010experiential and recent work showing LLMs improve via self-refinement zhao2024expel; gao2024self.
@@ 4 Learning from Instructor (LfI)
We investigate Learning from Instructor (LfI), where models acquire task knowledge via structured guidance, such as demonstrations, explanations, or feedback. We evaluate this setting across two dimensions: (1) Passive Consumption vs. Interactive Clarification, and (2) Scaling Learner.
@@
To examine the impact of interactivity in instruction-based learning, we adopt the MagpieMath (xu2024magpie) dataset, a high-quality math dataset generated autoregressively by Qwen2.5-Math-72B. An instructor model is first trained on this dataset and subsequently used to teach a separate learner model under two learning paradigms: (1) Passive Consumption, in which the learner receives only direct solutions from the instructor without any further interaction, and (2) Interactive Clarification, where the learner is allowed to ask clarification questions following each instructor response, and the instructor provides targeted, follow-up explanations. In this setting, the learner is restricted to a single clarification question per sample to ensure consistency across training examples. Both the initial answers and the clarification responses are aggregated as supervision for the learner. We experiment with four representative LLM families: LLaMA3.1-8B, Qwen2.5-7B, Mistral-7B, and Phi-3-mini—each used as both instructor and learner to test robustness across pairings. Learners are trained on outputs from matched instructors within each setting. Evaluation is conducted on eight diverse mathematical benchmarks: GSM8K (cobbe2021training), SVAMP (patel2021nlp), MATH (hendrycks2021measuring), NumGLUE (mishra2022numglue), SimulEq (koncel2016mawps), AQuA (ling2017program), SAT (zhong2023agieval), and MMLU-Math (hendryckstest2021), covering competencies from basic arithmetic to multi-step symbolic reasoning and standardized test preparation. Implementation and generation details are provided in Appendix A.3.
@@
As shown in Figure 2, learners trained under the Interactive Clarification paradigm consistently outperform those trained via Passive Consumption across all eight evaluation benchmarks. This performance gain highlights that LLMs are capable of leveraging interactive feedback to improve task understanding, resembling human-like active learning behaviors. Notably, the gains vary across model families. Mid-sized models such as Mistral-7B and Qwen2.5-7B show substantial improvements, suggesting a strong capacity to benefit from additional instructional signals. In contrast, the smallest model, Phi-3-mini, shows only marginal improvement, indicating limited ability to engage in or benefit from interactive clarification, which suggests that active learning capabilities are not uniform across models and may depend on model capacity or architecture.
Figure 3: Scaling learner models (Qwen2.5 variants) with a fixed Qwen2.5-72B instructor. Larger learners learn more effectively, and interactive clarification further improves outcomes.
@@ 4.2 Scaling Learner
4.3 Scale Instructor
Experiment Setup.
Main Result.
However, this trend breaks when the instructor model is extremely small (e.g., Qwen2.5-1.5B). In such cases, interactive clarification may introduce noise or misleading guidance, leading to degraded performance compared to passive learning. For example, on tasks such as GSM8K, MATH, and SAT in Figure 4, the interactive setting underperforms the passive counterpart, suggesting that low-quality instructors are unable to generate helpful responses to clarification queries, which in turn hinders learner improvement. As shown in Figure 4, larger instructor models consistently lead to better learning outcomes. When the instructor is strong (e.g., Qwen2.5-32B or 72B), interactive supervision significantly improves learner performance compared to passive instruction. This suggests that powerful instructors can provide useful clarifications that facilitate deeper understanding and generalization. However, when the instructor is too small (e.g., Qwen2.5-1.5B), the benefits of interactivity diminish or even become negative. In tasks such as GSM8K, MATH, and SAT, we observe that learners trained with Interactive Clarification underperform compared to those trained passively. This indicates that weak instructors may introduce noise or incorrect signals during interaction, ultimately hindering learning. These findings highlight the importance of instructor quality in interactive learning and suggest that poorly performing instructors can negatively affect the training process.
@@ 5.1 Structured Knowledge Injection in Competitive Environments
We evaluate whether conceptual knowledge improves strategic performance in multi-agent settings using TextArena (guertler2025textarena), a suite of competitive environments with symbolic rules and multi-turn dynamics. Each game involves two players, Player-0 and Player-1. We fix Player-0 as Qwen2.5-32B and vary Player-1 across four scales: Qwen2.5-1.5B, 7B, 14B, and 32B. Prior to gameplay, Player-1 either receives no guidance (without Concept) or is given natural language descriptions of game rules and strategies (with Concept), generated by Qwen2.5-32B. For each environment and model pair, we run 20 matches and report Player-1’s win rate averaged over these games. Environments include Checkers (CH), Poker (PK), Stratego (ST), Tic Tac Toe (TT), Truth and Deception (TD), and Ultimate Tic Tac Toe (UTT). Performance is measured as Player-1’s win rate against the fixed opponent. Results are shown in Figure 5. Implementation details are provided in Appendix A.4.
@@
Table 1: Accuracy on six tasks with and without concept input (baseline / +concept). Higher scores indicate better subgoal completion or reasoning accuracy. Darker cell colors indicate larger improvements from concept input; white means no change.
Qwen 2.5 7b 14b 32b 72b
LogicGrid 0.36/0.42 0.51/0.54 0.53/0.56 0.54/0.57
NaLogic 0.13/0.16 0.14/0.18 0.17/0.20 0.21/0.23
Plan 0.14/0.17 0.32/0.34 0.33/0.34 0.33/0.35
AlfWorld 0.27/0.29 0.30/0.32 0.33/0.34 0.33/0.33
ScienceWorld 0.16/0.19 0.26/0.29 0.33/0.35 0.41/0.42
BabyAI 0.40/0.42 0.41/0.44 0.46/0.46 0.51/0.51
@@ 5.2 Conceptual Generalization in Logic and Planning Tasks
To evaluate rule-based generalization, we include six tasks centered on explicit symbolic or logical structures. LogicGrid and NaLogic feature structured and narrative logic puzzles (mandell1986fantastic; dudeney1907canterbury), testing deduction under constraint. Plan (vallati20152014) translates classical symbolic planning domains (e.g., Gripper, Barman) into natural language settings to assess structured action reasoning. Additionally, tasks from AlfWorld (shridhar2020alfworld), ScienceWorld (wang2022scienceworld), and BabyAI (chevalier2018babyai), models must complete goal-oriented tasks requiring commonsense reasoning, procedural knowledge, and spatial navigation. For the concept-injected setting, we use Qwen2.5-72B to generate static conceptual hints, which are provided as auxiliary input to smaller models during evaluation. Results are presented in Table 1 and Figure 5. We use subgoal-based annotations to enable fine-grained progress tracking. Implementation details are provided in Appendix A.4.